The intensity distribution of solar radiation peaks at a wavelength of approximately 500 nm. Assuming the Sun radiates like a black body, calculate its surface temperature in K.


1
Expert's answer
2020-08-28T06:25:10-0400

λpeak = 500 nm = 5*10-7 m

According to Wien's displacement law,

λpeak = b/T, where b is constant equal to 2.898*10-3 m*K

Therefore, T = b/λpeak = 2.898*10-3 / 5*10-7 = 5.796*103 K, or 5796 K

Answer: 5796 K